18:00: we start setting up the system: we acquire new Dark: 20141129_020106

TN Notes
20141129_020106 w/o light
20141129_020618
20141129_020759

18:18: we point Hip117974 on FLAO, bin2, 400Hz, 70phot/subap

We start laser propagation but the beacons are really out, we acquire with MATS. On top of this we have problems with collimation.

!8:40: FLAO at 550Hz, ARGOS at 510Hz.

PBL with ASM: shell panic probabli due to AARB restart

LAN focus at -75k steps, LAS ToF at 85us.

TN Notes
20141129_030811 CL just on FLAO, G=1.5,0.6,0.4
20141129_031131 ", and JCL, Gj=0.1
20141129_041324 CL FLAO+ARGOS, G=0.4,0.3,0.1
20141129_041406
20141129_041610 G=0.4,0.4,0.3

20:18: we move the periscope to better recenter the 3 LGS on LGSW FoV

We define a new PatCam target: 20141130_013400

PatCam X Y
Blue 178 125
Yellow 181 121
Red 180 131

After that we acquire a new PatCam to Periscope IM and REC: 20141129_050500

We set as new laser acquisition target: 20141129_052400

21:35: we preapare the REC for a single LGS having TCoG48 slope computation algorithm, we use the Yellow LGS:

COMB. REC TAG REC LGSW FLAO REC COMB. REC
Yellow TCoG48 only 20141129_054100 - 20141129_054200
Yellow TCoG48 BIN2 20141129_054100 20130916_221257 20141129_054300

22:38: we recollimate (it takes a while!)

we close the loop just on FLAO, 150 modes, OK

we close on ARGOS + FLAO, 500Hz

TN Notes
20141129_071327 G=0.2,0.1,0.1
20141129_071633 G=0.3,0.2,0.2

We acquire the IM between Periscope and LGSW slopes ad we close the loop on it to prevent the LGS drift

Periscope LGSW TT REC: 20141129_072736

TN Notes
20141129_073330 G=0.3,0.2,0.2
20141129_073446

23:39: we put a star on FLAO and we try to close the TT loop on that

TN Notes
20141129_075351 G=0.5,0.4,0.3

we set the Z stage on FLAO to 27mm, the best position for LUCI (when the telescope is collimated we have the stage at 42mm instead).

We optimize the astigmatism corrector angle: 800 steps

We optimize the ToF: 85us are fine for Z stage at 42mm, we go with the stage at 27mm and we measure 10um of focus on the TS. We close the TS, the focus term is offloaded on the LGSW signals. We reoptimize the ToF: 85.8us.

TN Notes
20141129_094113 G=0.4,0.3,0.2
20141129_094206
20141129_094530 G=1.2,0.4,0.3
20141129_094546
20141129_094734 G=1.2,0.6,0.5
20141129_094832
20141129_095054 G=1.5,0.6,0.5
20141129_095152
20141129_095248 LUCI image #42
20141129_095456 G=2.2,0.6,0.5
20141129_095626
20141129_095918 G=2.2,0.7,0.6
20141129_095953
20141129_100226 OL
20141129_100257
20141129_100327

2:00: we set the speed to 700Hz but everything dies.

HIP 208211

LUCI images, from SR notebook. Date synchronization is unknown.

See the table below and raw_20141128.txt for a list of LUCI images.

Date LUCI Band FWHM Notes
20141129_111800 20141128-46 ?    
  20141128-47 ?    
  20141128-48 ? 0.375 Loop on LGS high gain
  20141128-49 ? 0.75 Loop paused
  20141128-50 ?   Truth stopped
20141129_113000 20141128-51 ? 0.25 Truth stopped
20141129_113300 20141128-52 ? 0.625 Loop paused
20141129_114100 20141128-53 ?   Loop on Yellow LGS only

2:45: we have repointed the target and we close again

TN Notes
20141129_112100 G=1.0,0.1,0.05
20141129_112158 G=1.5,0.2,0.1
20141129_112245 G=2.0,0.4,0.2
20141129_112400 G=2.2,0.7,0.6
20141129_112449 G=2.2,0.6,0.5
20141129_112528 G=2.2,0.6,0.5, seeing=1.2"
20141129_112651 OL
20141129_112729
20141129_112903 G=2.2,0.6,0.5, no TS
20141129_112929 G=2.2,0.6,0.5
20141129_113214 G=2.2,0.6,0.5, LUCI image # 50

Slope_offset for TS: 20141129_113400

TN Notes
20141129_113559 OL
20141129_113725 G=2.2,0.6,0.5

3:40: we open, load the REC for single LGS and close again

TN Notes
20141129_114334 G=2.2,0.6,0.5
20141129_114413
20141129_114445
20141129_115103 G=2.2,0.4,0.3, TS on
20141129_115136
20141129_115233 G=2.2,0.5,0.4

NGC 2419

This is the cluster of the image used for the press release (see SR's email on 141202 "Argos GLAO closed loop").

See the table below and raw_20141128.txt for a list of LUCI images.

Date LUCI Band FWHM Notes
20141129_125900 20141128 57-70 H    
20141129_130600 20141128 71-90 K    
20141129_130900 20141128 91-100 Ks   some (all?) Open Loop
20141129_131400 20141128 101-120 J    
20141129_131600 20141128 121-140 Ks    
20141129_131800 20141128 141-150 J   Open Loop
20141129_131900 20141128 151-160 H   Open Loop
20141129_132800 20141128 161-200 H   Closed Loop
20141129_134300 20141128 201-210 H   Sky
  20141128 211-220 Ks   Sky
  20141128 221-270 J   Sky
  20141128 271-300 H   Sky
  20141128 301-330 Ks   Sky

Back on 3 LGS:

TN Notes
20141129_130416 G=3.0,0.6,0.5
20141129_130555
20141129_131608 G=2.0,0.4,0.2, seeing=*1.2"*
20141129_131915
20141129_132939
20141129_133220 G=2.0,0.5,0.3, PC opening time=2us
20141129_133257
20141129_133512 G=2.0,0.6,0.5

-- MarcoBonaglia - 17 Dec 2014
I Attachment Action Size Date Who Comment
raw_20141128.txttxt raw_20141128.txt manage 20 K 13 Jan 2015 - 16:09 LorenzoBusoni List of LUCI2 images taken on 2014 11 28
Topic revision: r4 - 14 Jan 2015, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback